1. Assuming that the pure rate of interest is 2%, and investors require an inflation premium of 3.5% and a risk premium of 6% to invest in a certain security, calculate the following rates using the multiplicative form of the Fisher model:


The nominal rate of interest on the security


The real rate of interest on the security


The risk-free rate of interest on securities of this maturity 


  (2 decimal places)


 


2.  An Inyo County California municipal bond is currently yielding 4.2%. What after-tax yield would you receive if you are in the following circumstances:


You are in a 28% federal tax bracket, and, as a California resident, you are in the 5% state tax bracket.


You are in a 33% federal tax bracket, and, as a Utah resident, you pay a 4% state income tax.


You are in a 15% federal tax bracket, and, as a Nevada resident, you pay no state income tax because Nevada has no income tax. 


 ( 2 decimal places.)


 


3. The expected return on a share of ExxonMobil stock in the U.S. is 15.6% while the expected return on a share of Royal Dutch Shell stock is 12.6% in the Netherlands. If the pure rate of return is 2% in both countries and the required risk premium is 6% for each company's stock, what is the long-term expected inflation rate in each country if the multiplicative form of the Fisher model is used in making the calculations?  


 (2 decimal places.) 
Answer #1
Part 1Risk free rate of interest = ((1+pure interest) (1+inflation premium))-1



Pure interest = 2%

Inflation Premium = 3.5%



Risk free rate of interest = ((1+2%)(1+3.5%))-1

Risk free rate of interest = ((1.02)(1.035))-1

Risk free rate of interest = 5.57%



Nominal interest rate is the pure interest rate + inflation premium

Nominal interest rate = 2% + 3.5%

Nominal interest rate = 5.5%



Fisher equation

(1+nominal interest rate)= (1+ real interest rate) (1+inflation rate)



Real interest rate = (1+nominal interest rate)/(1+inflation rate)-1

Real interest rate = (1+nominal interest rate)/(1+inflation rate)-1

Real interest rate = ((1+5.5%)/(1+3.5%))-1

Real interest rate = (1.055/1.035)-1

Real interest rate = 1.93%
Part 2Bond Current Yield = 4.2%


a)Federal tax = 28%

State tax=5%



Effective tax rate = 5%*(1-28%)+28%

Effective tax rate = 3.6%+28%

Effective tax rate = 31.6%



After tax yield = Bond yield/(1-effective tax rate)

After tax yield = 4.2/(1-31.6%)

After tax yield = 4.2/(1-31.6%)

After tax yield = 4.2/(1-31.6%)

After tax yield = 4.2/0.684

After tax yield = 6.14%


b)Utah resident

In this case we cant deduct the state tax payment in federal tax return



After tax yield = Bond yield/(1-Federal tax rate- state tax rate)

After tax yield = 4.2/(1-33%-4%)

After tax yield = 4.2/0.63

After tax yield = 6.67%


c)There is only federal tax



After tax yield = Bond yield/(1-Federal tax rate)

After tax yield = 4.2/(1-15%)

After tax yield = 4.2/0.85

After tax yield = 4.9%


answered by: gavin